Find the average value gave of the function g on the given interval. g(x) = 5Vx. [1,27]
Answer #1

The average value of y = f(x) on [a b] is tang = 1 fox dx b-a. goni = 5. dx = 5. (2)3 on [1, 27]. have - Jon - sa cont de 5 FThank you!
Answer #2

OLUTION :


Average value (gave)  of the given function g(x) in interval (1, 27)


   27

= ∫    g(x) dx  / (27 - 1)

   1


              27

= 1/26   ∫    5 (x^(1/3)  dx

              1

                                 27

= 5/26* 3/4 [ x^4/3) ]                            

                                 1                                                       


= 15/104 * [27^(4/3) - 1^(4/3)]

= 15/104 * 80

= 11.54 (ANSWER)

answered by: Tulsiram Garg